# Topological flow data analysis for transient flow patterns: a graph-based approach

### v1

@@ 3.2.2. Information transfer between corner regions
We here employ Convergent Cross Mapping (CCM) developed by Sugihara et al. (2012) to measure causality among flow pattern changes. This causal inference can be classified as observational, following the classification used for Markov processes in Baldovin et al. (2020). This approach investigates if knowledge of one variable is useful in predicting future values of another. While interventional methods are commonly employed in fluid mechanics owing to their direct relevance to active flow control, they are unsuitable for our present analysis based on the qualitative geometric information extracted by TFDA.
@@
To encode the instantaneous flow state based on COT, the state $v^{i}_{k}$ of each corner $i=0,1,2$ at time $t=t_{k}$ is defined as follows: $v^{i}_{k}=3$ if $b_{++}$ is present inside the parenthesis of $c_{+}^{i}$; otherwise, it is the smaller of the number of $c_{+}^{i}$ and $2$. We form a state vector at the $i$-th corner using time-delayed embedding:
@@
with ${\mathcal{E}}=30$ and $\tau=30$. These values of ${\mathcal{E}}$ and $\tau$ are determined by maximising the accuracy of the autoregression. Figure 14 shows the evolution of the state vectors at the three corners by plotting the first principal components of $u^{i}_{k}\ (k=0,\ldots,5000)$ for $i=1,2,3$ on the $y$-axis against time on the $x$-axis. Note that the unit time is $0.01$. For $Re=14000$ (Figure 14(a)) and $Re=15250$ (Figure 14(b)), clear periodicity is observed, while no apparent pattern can be seen for $Re=15400$ (Figure 14(c)) and $Re=16000$ (Figure 14(d)).
@@
To quantify spatial and temporal information transfer between corners, we quantify how accurately one local structure’s state vector predicts another’s. Since $c^{2}_{+}$ remains static for low Reynolds numbers, we focus on the states of $c^{0}_{+}$ and $c^{1}_{+}$. Using the time-delayed embedding $u^{i}_{k}$ defined earlier, CCM evaluates prediction accuracy through correlation coefficients, using increasingly larger historical datasets until convergence. In a deterministic system, the dynamics of the driver’s variables is embedded within the effect variables’ dynamics. Higher prediction accuracy of one variable using another’s state therefore indicates that the predicted variable causally influences the predictor variable. In Figure 15, we plot two key relationships. The orange curve (“c1:c0”) shows the correlation between the predicted and actual state vectors at $c^{0}_{+}$, where predictions use $c^{1}_{+}$’s state vector. A high correlation in this case reveals that $c^{0}_{+}$’s state can be reconstructed from $c^{1}_{+}$’s dynamics, indicating that $c^{0}_{+}$ causally drives $c^{1}_{+}$. The blue curve (“c0:c1”) represents the reverse relationship. Up to $Re=15250$, both curves maintain values near $1$. It indicates an indistinguishable bidirectional causality, reflecting the periodic evolution of streamline patterns observed in Section 3.2.1. However, at approximately $Re=15400$, causation from $c^{1}_{+}$ to $c^{0}_{+}$ (blue) weakens, while causation from $c^{0}_{+}$ to $c^{1}_{+}$ (orange) persists up to $Re=16000$. This suggests that as Reynolds number increases, topological pattern changes at the top left corner increasingly drive changes at the bottom left corner. Notably, this causality difference between periodic and chaotic regimes, identified through our observational approach based on COT, is not captured by interventional linear response function analysis as discussed in Appendix A.
Figure 15. Convergent cross mapping (CCM) analysis of the corner states. The correlation coefficient between true and predicted states (the $y$-axis) is plotted against Reynolds number (the $x$-axis). The blue curve shows prediction accuracy when using $c^{0}_{+}$ to predict $c^{1}_{+}$, while the orange curve shows the reverse relationship. Higher correlation coefficients indicate a stronger causal influence from the predicted state to the predictor state, following CCM theory.
@@ 4. Summary
TFDA is a new method of topological data analysis for two-dimensional vector fields. It allows us to extract all local structures to be identified geometrically and to describe their global structure as a tree (COT) and its string representation (COT representation) based on the theorems by Poincaré–Hopf and Poincaré–Bendixson. Each node label (COT symbol) of the tree expresses a local orbit structure in the flow. The global configuration among local orbit structures are expressed as edges of the tree without ambiguity. Since the snapshots of the flow patterns are converted into a series of COTs or COT representations, the continuous evolution of two-dimensional incompressible flows is reduced to a discrete dynamical system between topologically equivalent flow patterns.
@@
We apply TFDA to the evolution of the lid-driven cavity flow for the Reynolds number ranging from $Re=14000$ to $Re=16000$, which has been well investigated as a test bed for many existing analysis methods. When the Reynolds number is $Re=14000$, five topologically different streamline patterns are observed, and their corresponding COT representations repeat periodically throughout the evolution. By examining the changes in the COT representations of the flow patterns, we can theoretically identify the marginal flow pattern between them, even if it is not explicitly contained in the simulation data. In addition, by tracking the change of the COT symbol $c_{+}$ contained in the COT representations, which represent the existence of a recirculating flow domain at a cavity corner, we can express the evolution of the flow as a cyclic transition graph. Suppose that the dynamics is regarded as a Markov process on the transition graph. Then the period-five dynamics of the flow patterns are extracted as a pentagonal distribution of the eigenvalues of the transition matrix on the unit circle. This is a remarkable feature of TFDA, which describes transient pattern dynamics using the graph-theoretic approach. We apply the same analysis to the flow evolutions in the cavity from $Re=14000$ to $Re=16000$. We find that the dynamics of the flow change from periodic to chaotic at around $Re=15400$, which agrees with the critical Reynolds number confirmed by other existing analysis methods. It suggests that TFDA captures the qualitative change in the behaviour of these dynamics.
To complement our graph analysis, we employ convergence cross mapping (CCM) to investigate causal relationships between topological flow pattern changes across different cavity corners. This method analyses vectorised information from the COT representations at each corner to determine how changes in one corner’s recirculating domain influence another. In the periodic flow regime ($Re=14000-15300$), we detect no significant causal relationships between pattern changes at different corners. However, in the chaotic regime ($Re=15500-16000$), we observe asymmetric causality: changes in the top left corner’s topological pattern exhibit a stronger causal influence on the bottom left corner than vice versa.
@@
TFDA’s primary strengths derive from two key characteristics: robustness and interpretability. Unlike traditional methods such as POD, which are sensitive to measurement noise and numerical artefacts, COTs maintain their structural integrity under perturbations owing to their topological nature. This characteristic makes the analysis particularly reliable for experimental data where noise is inevitable. A notable example of TFDA’s robustness is its application to left ventricular flows (Sakajo and Itatani, 2023), where TFDA successfully identified vortex structures even in noisy clinical flow data obtained through echocardiography. The interpretability of COTs stems from their direct connection to physically meaningful flow structures. Each symbol in a COT corresponds to a specific qualitative feature in the flow, enabling immediate physical interpretation. This attribute contrasts with POD, where the physical meaning of individual modes occasionally remains obscure.
@@
On the other hand, however, the COT methodology faces important limitations, particularly in its application to three-dimensional flows. The current formulation relies inherently on two-dimensional topology, necessitating the projection of three-dimensional flows onto appropriate two-dimensional sections. While this approach has proven effective in specific applications, including our analysis of left ventricular flow patterns (Sakajo and Itatani, 2023), it requires careful selection of projection planes to capture the most relevant flow features. This dimensional restriction, though limiting, often aligns well with practical applications where key flow features develop primarily in specific planes or where experimental measurements are naturally restricted to particular cross-sections. Future research directions may include extending the COT framework to a fully three-dimensional topological analysis.
@@ Appendix A Linear response function analysis
To highlight the differences between observational and interventional approaches in detecting causality, we present a numerical analysis of the linear response function under spatially localised forcing. This analysis provides complementary insights to the convergent cross mapping results discussed in Section 3.2.2. Let us recall the formalism of the linear response function (Marconi et al., 2008). In the numerical simulation of the lid-driven cavity, the flow field reaches a statistically steady state such as a time-periodic state or a chaotic state at $t=T_{s}$. Then we start to add an external forcing $f_{\omega}({\bm{x}},t)$ to the vorticity equation at this time $T_{s}$ and switch it off at $t=T_{s}+T_{f}$. The forcing is localized in space and the duration $T_{f}$ is suitably chosen, as we will discuss later. We denote the velocity and the vorticity with the forcing by $\widetilde{{\bm{u}}}$ and $\widetilde{\omega}$. The vorticity $\widetilde{\omega}$ obeys the forced equation,
@@
We numerically solve both the forced equation (10) and the unforced vorticity equation. The velocity and the vorticity without the forcing are denoted by ${\bm{u}}$ and $\omega$. If the forcing is small enough, the difference of the two vorticity fields can be expressed with the linear response function $G_{\omega}({\bm{x}},t|{\bm{x}}^{\prime},t^{\prime})$ as
@@
Now, for a moment, let us suppose formally that the forcing is localized at ${\bm{x}}_{f}$ and $T_{s}$, that is, $f_{\omega}({\bm{x}},t)=a\delta({\bm{x}}-{\bm{x}}_{f})\delta(t-T_{s})$, where $\delta(x)$ is the Dirac delta function and $a$ is the amplitude of the forcing. Then we have
@@
Here, the width of the forcing is set to $\sigma=1/8$. We set ${\bm{x}}_{f}$ either to the centre of $D_{0}$ ($(x_{f},y_{f})=(-3/4,3/4)$), or to the centre of $D_{1}$ ($(x_{f},y_{f})=(-3/4,-3/4)$). The former case is denoted by ${\bm{x}}_{f}\in D_{0}$ and the latter by ${\bm{x}}_{f}\in D_{1}$. This forcing added in the $D_{1}$ region and the corresponding vorticity forcing are shown in Figure 16. Both forcings are localized. The amplitude of the forcing is set to $a=\langle\psi\rangle_{D_{j}}\epsilon/\Delta t$, where $\Delta t$ is the time step of the simulation ($\Delta t=10^{-3}$ for $Re\leq 16000$) and $\langle\psi\rangle_{D_{j}}$ is the space-time average of the streamfuntion without the forcing in the region $D_{j}$. We set the parameter to $\epsilon=1.0\times 10^{-7}$ and check that smaller values of $\epsilon$ than this value give the same result. The duration of the forcing is set to $T_{f}=10$, which is about the same as the maximum delay time $910\Delta t\times\mathcal{E}\times\tau=$ 9 used for the embedding in Section 3.2.2 (recall that the COT representations are obtained at every 10 steps). We also set $T_{f}$ to smaller values, e.g., $T_{f}=0.5$, and check that the qualitative results do not change.
In the following we observe $a^{-1}\Delta\omega$ to check the relation (14). More specifically, we consider the average of the difference over the region $D_{1}$ by adding the forcing in the region $D_{0}$, which is denoted by
@@
Similarly we define $\left\langle a^{-1}\Delta\omega\right\rangle_{D_{1}\to D_{0}}$. The average can be regarded as a suitable average of the linear response function $G_{\omega}({\bm{x}},t|{\bm{x}}^{\prime},t^{\prime})$ in view of equation (11). We also calculate the difference of the streamfunction in the same manner (we only consider the scalar quantities here).
@@
For $Re=16000$, we plot the averages of the vorticity and streamfunction differences as a function of time in Figure 17. The differences are normalised by the forcing amplitude and the standard deviations of the observed regions. The amplitudes of the vorticity differences do not agree with our anticipation (13) or (14) as shown in Figure 17(a). On the other hand, for the streamfunction, with which we identified the structures $c^{1}_{+}$ and $c^{0}_{+}$, the amplitude of the difference for $D_{1}\to D_{0}$ is non-zero in $t\leq 15$, while the amplitude for $D_{0}\to D_{1}$ is around zero, as observed in Figure 17(b). However, in $t-t_{f}>15$ the amplitudes become comparable. From this behaviour, we infer that $c^{1}_{+}$ has a causal influence on $c^{0}_{+}$ up to around $t-t_{f}=T_{f}$ (Baldovin et al., 2020) and that at later times no difference in causality is detected. Therefore, it is not straightforward to conclude which corner eddy is causal on another one by observing the difference of the amplitudes. For the lower Reynolds number $Re=14000$, where the flow is periodic in time, we plot the averaged differences as the linear response functions in Figure 18. Their overall shapes remain similar to those for $Re=16000$. Hence, we hardly find qualitative differences in linear responses across these Reynolds numbers. This may be owing to the linearity of their governing equations and comparable Péclet numbers despite the underlying flow transitioning from periodic to chaotic states. As a result, we find virtually no difference in sensitivity for periodic and chaotic evolutions. In this sense, the observational CCM approach based on TFDA can clarify the geometric causality between topological changes in flow patterns that the standard sensitivity analysis cannot capture. The discrepancy between CCM and the linear response analyses can be attributed to their fundamentally different approaches to causality detection. CCM operates in the framework of state-space reconstruction, utilising Taken’s embedding theorem to predict future states from historical time series data. In the periodic regime illustrated in Figure 10(a), the COT representations exhibit highly regular patterns. Under these conditions, the system’s future states can be predicted from any instantaneous state with equal accuracy, regardless of which corner’s information is used. This makes causal relationships appear symmetric in the CCM analysis. Linear response analysis, in contrast, provides a mechanistic view of causality by examining the system’s response to controlled perturbations. This approach directly probes the spatio-temporal propagation of disturbances through the flow field. This methodological distinction highlights the complementary nature of observational (CCM) and interventional (linear response) approaches in understanding complex fluid systems.
@@
Finally, we discuss the relations of our work to recent studies on causality in fluid mechanics. In three-dimensional turbulent flows, a standard way is to apply POD first and then study causal relationships between the POD modes. In Lopez-Doriga et al. (2024), the square duct flow was studied to uncover the cause of the secondary flow with the POD and the Granger causality analysis. In Martínez-Sánchez et al. (2023), the wake behind a circular cylinder mounted on a wall was analysed with the POD and the entropy-transfer method for the causality. A time-series-based analysis for the sub-grid energy fluxes with different grid scales in homogeneous and isotropic turbulence was performed in Lozano-Durán and Arranz (2022) with the entropy-transfer method. The most elaborated interventional study to date in this field is Lozano-Durán et al. (2021) where the wall-bounded, minimal-flow-unit turbulence was analysed by numerically removing certain unstable modes to find the primary cause of the sustainment of the turbulent fluctuations by the mean flow. In this paper, we studied the causality between the corner eddies represented by the COT symbols $c^{0}_{+}$ and $c^{1}_{+}$ as discrete modes in the lid-driven cavity flow. The difference from the previous works is that we used new modes and applied both approaches of the causal analysis. The advantages of the new modes, COTs, are the immediate interpretability and the robustness to noises owing to the topological approach. The COTs are not much influenced by noise, unlike POD. The disadvantage is that it does not apply directly to three-dimensional flows, but we need to consider the projection of the flow onto an appropriate two-dimensional section as we discussed in Summary.

### v2

@@ 4.5. Intrinsic causality analysis between flow patterns in corner regions
Here we employ Convergent Cross Mapping (CCM) developed by Sugihara et al. (2012) to measure causality among flow pattern changes. This causal inference can be classified as observational, following the classification in Baldovin et al. (2020). This approach investigates whether knowledge of one variable is useful in predicting future values of another. This framework is well suited to analysing sequences of COT representations. An interventional analysis based on continuous quantities, which is commonly employed in fluid mechanics because of their direct relevance to active flow control, is presented in Appendix A.
@@
To encode the instantaneous flow state based on COT symbols, the state $v^{i}_{k}$ of each corner $i=0,1,2$ at time $t=t_{k}$ is defined as follows: $v^{i}_{k}=3$ if $b_{++}$ is present inside the parentheses of $c_{+}^{i}$; otherwise, it is the smaller of the numbers of $c_{+}^{i}$ and $2$. We form a state vector at the $i$-th corner using time-delayed embedding:
@@
with ${\mathcal{E}}=30$ and $\tau=30$. These values of $\mathcal{E}$} and $\tau$ are determined by maximising the accuracy of the autoregression. Note that the unit time is $0.01$. For $Re=14000$ (Figure 21(a)) and $Re=15250$ (Figure 21(b)), clear periodicity is observed, while no apparent pattern can be seen for $Re=15400$ (Figure 21(c)) and $Re=16000$ (Figure 21(d)).
@@
To quantify the spatial and temporal causal relation between corners, we quantify how accurately one local structure’s state vector predicts another’s. Since $c^{2}_{+}$ remains static for low Reynolds numbers, we focus on the states of $c^{0}_{+}$ and $c^{1}_{+}$. Using the time-delayed embedding $u^{i}_{k}$ defined earlier, CCM evaluates the precision of the prediction through correlation coefficients, using increasingly larger historical datasets until convergence. In a deterministic system, the dynamics of the driver’s variables is embedded within the effect variables’ dynamics. The higher prediction accuracy of one variable using the state of another therefore indicates that the predicted variable causally influences the predictor variable. In Figure 22, we plot two key relationships. The orange curve (“c1:c0”) shows the correlation between the predicted and actual state vectors at $c^{0}_{+}$, where the predictions use the state vector of $c^{1}_{+}$. A high correlation in this case reveals that the state of $c^{0}_{+}$ can be reconstructed from the dynamics of $c^{1}_{+}$, indicating that $c^{0}_{+}$ causally drives $c^{1}_{+}$. The blue curve (“c0:c1”) represents the reverse relationship.
Up to $Re=15250$, both curves maintain values close to $1$, indicating an indistinguishable bidirectional causal relationship. This reflects the periodic evolution of the streamline patterns described in Section 4.2, where the topological changes at different corners occur in a coordinated manner. However, around $Re=15400$, the causation from $c^{1}_{+}$ to $c^{0}_{+}$ (blue) weakens, while the causation from $c^{0}_{+}$ to $c^{1}_{+}$ (orange) persists up to $Re=16000$. This asymmetry suggests that, in the aperiodic regime, topological pattern changes at the top left corner increasingly drive changes at the bottom left corner. In other words, the transition to aperiodic behaviour is accompanied by a reorganisation of the interaction structure among corner eddies.
These results provide a picture of the transition observed in the COT dynamics. In the periodic regime, the corner structures evolve in a mutually coupled manner, producing synchronous pattern transitions. In contrast, once the flow becomes aperiodic, the influence among corners becomes uneven, and the evolution of the system is effectively initiated by structural changes in the top left corner. Thus, causal analysis based on the COT representation reveals how the hierarchy of interactions among coherent flow structures changes across the transition. In Appendix A, to complement the present observational causal analysis, we further investigate the relation between the two corners through an interventional analysis, which is more closely aligned with control-oriented studies in fluid mechanics.
Figure 22. Convergent cross mapping (CCM) analysis of the corner states. The correlation coefficient between true and predicted states (the $y$-axis) is plotted against Reynolds number (the $x$-axis). The blue curve shows prediction accuracy when using $c^{0}_{+}$ to predict $c^{1}_{+}$, while the orange curve shows the reverse relationship. Higher prediction accuracy (higher correlation between the prediction and the truth) indicate a stronger causal influence from the predicted state to the predictor state, following CCM theory.
@@ 5. Summary
TFDA is a new method of topological data analysis for two-dimensional vector fields. It allows us to extract all local structures to be identified geometrically and to describe their global structure as a tree (COT) and its string representation (COT representation). Each node label (COT symbol) of the tree expresses a local orbit structure in the flow. The global configuration among local orbit structures is expressed as edges of the tree without ambiguity. Since the snapshots of the flow patterns are converted into a series of COT representations, the continuous evolution of two-dimensional incompressible flows is reduced to a discrete dynamical system between topologically equivalent flow patterns.
@@
We apply TFDA to the evolution of the lid-driven cavity flow for the Reynolds number ranging from $Re=14000$ to $16000$, which has been well investigated as a test bed for many existing analysis methods. When the Reynolds number is $Re=14000$, six topologically different streamline patterns, called topological states, are observed, and their corresponding COT representations repeat periodically throughout the evolution. By examining the changes in the COT representations of the flow patterns, we can theoretically identify the marginal flow pattern between them, even if it is not explicitly contained in the simulation data. In addition, by tracking the time series of the COT representations, we can express the evolution of the flow as a graph that describes a cyclic transition between topological states.
Using TFDA for the time series of COT representations for various Reynolds numbers, we demonstrate how one can understand the physics of the lid-driven cavity flow. We have conducted the following five analyses. First, we argue that the transition in flow complexity can be studied by the number of topological states identified with TFDA. We observe critical behaviour when the Reynolds number exceeds $Re=15400$. Quantitative aspects, such as the value of the critical exponent, are affected by how one samples the Reynolds number, the time series of the streamfunction, and the threshold of TFDA to calculate COT. Second, we show that the period of periodic flow dynamics can be estimated by the time series of COT representations. Third, we argue that the variation of physical quantities of interest, such as energy and enstrophy, can be interpreted in terms of local flow patterns of topological states and the sequence thereof. This can be used in addition to classifying or predicting a flow state. In particular, here, we focus on local maxima and minima of enstrophy and the corresponding topological states. Fourth, we perform TFDA for flows of various Reynolds numbers, and assign serial numbers (IDs) to the topological states in the order of their appearance. We then investigate their distribution and the rate of occurrence. Consequently, we found that topological states with simple structures that appear in periodic motion at low Reynolds numbers remain observed for a long time in chaotic motion at high Reynolds numbers. Finally, we employ convergence cross mapping (CCM) to investigate causal relationships between topological flow pattern changes across different cavity corners. This method analyses vectorised information from the COT representations at each corner to determine how changes in one corner’s recirculating domain influence another. In the periodic flow regime (Re=1400015300Re=14000-15300), we do not detect significant causal relationships between pattern changes at different corners. In the chaotic regime (Re=1550016000Re=15500-16000), we observe asymmetric causality: changes in the topological pattern at the top left corner exhibit a stronger causal influence on the bottom left corner than vice versa.
@@
The primary strengths of TFDA derive from two key characteristics: robustness and interpretability. Unlike traditional methods such as POD, which are sensitive to measurement noise and numerical artifacts, COT maintain their structural integrity under perturbations owing to their topological nature. This characteristic makes the analysis particularly reliable for experimental data where noise is inevitable. A notable example of the robustness of TFDA is its application to left ventricular flows Sakajo and Itatani (2023), where TFDA successfully identified vortex structures even in noisy clinical flow data obtained by echocardiography. The interpretability of COT comes from their direct connection to physically meaningful flow structures. Each symbol in a COT corresponds to a specific qualitative feature of the flow, enabling immediate physical interpretation. This attribute contrasts with POD, where the physical meaning of individual modes occasionally remains obscure.
@@
However, on the other hand, the COT methodology faces important limitations. First, since the streamfunction is not Galilean invariant, it should be noticed that the results obtained here with TFDA are coordinate dependent. Second, the current formulation is inherently based on two-dimensional topology, necessitating the projection of three-dimensional flows onto appropriate two-dimensional sections. Although this approach has proven effective in specific applications, including our analysis of left ventricular flow patterns Sakajo and Itatani (2023), it requires careful selection of projection planes to capture the most relevant flow feature. This dimensional restriction, though limiting, often aligns well with practical applications where key flow features develop primarily in specific planes or where experimental measurements are naturally restricted to particular cross sections. Future research directions may include extending the COT framework to a fully three-dimensional topological analysis.
@@ Appendix A Linear response function analysis
To highlight the differences between observational and interventional approaches in detecting causality, we present a numerical analysis of the linear response function under spatially localised forcing. This analysis provides complementary insights to the convergent cross mapping results discussed in Section 4.5. Let us recall the formalism of the linear response function Marconi et al. (2008). In the numerical simulation of the lid-driven cavity, the flow field reaches a statistically steady state such as a time-periodic state or a chaotic state at $t=T_{s}$. Then we start adding an external forcing $f_{\omega}({\bm{x}},t)$ to the vorticity equation at this time $T_{s}$ and stop it at $t=T_{s}+T_{f}$. The forcing is localized in space and the duration $T_{f}$ is suitable, as we will discuss later. We denote the velocity and the vorticity with the forcing by $\widetilde{{\bm{u}}}$ and $\widetilde{\omega}$. The vorticity $\widetilde{\omega}$ obeys the forced equation,
@@
We numerically solve both the forced equation (11) and the unforced vorticity equation. The velocity and the vorticity without forcing are denoted by $\bm{u}$} and $\omega$. If the forcing is small enough, the difference between the two vorticity fields can be expressed with the linear response function $G_{\omega}({\bm{x}},t|{\bm{x}}^{\prime},t^{\prime})$ as
@@
Now, for a moment, suppose formally that the forcing is localized at ${\bm{x}}_{f}$ and $T_{s}$, that is, $f_{\omega}({\bm{x}},t)=a\delta({\bm{x}}-{\bm{x}}_{f})\delta(t-T_{s})$, where $\delta(x)$ is the Dirac delta function and $a$ is the forcing amplitude. Then we have
@@
Here, the width of the forcing is set to $\sigma=1/8$. We set ${\bm{x}}_{f}$ either to the centre of $D_{0}$ ($(x_{f},y_{f})=(-3/4,3/4)$), or to the centre of $D_{1}$ ($(x_{f},y_{f})=(-3/4,-3/4)$). The former case is denoted by ${\bm{x}}_{f}\in D_{0}$ and the latter by ${\bm{x}}_{f}\in D_{1}$. This force added in the $D_{1}$ region and the corresponding vorticity forcing are shown in Figure 23. Both forcing are localized. The forcing amplitude is set to $a=\langle\psi\rangle_{D_{j}}\epsilon/\Delta t$, where $\Delta t$ is the simulation time step ($\Delta t=10^{-3}$ for $Re\leq 16000$) and $\langle\psi\rangle_{D_{j}}$ is the space-time average of the flow function without forcing in the region $D_{j}$. We set the parameter to $\epsilon=1.0\times 10^{-7}$ and check that values smaller than $\epsilon$ than this value give the same result. The forcing duration is set to $T_{f}=10$, which is about the same as the maximum delay time $10\Delta t\times\mathcal{E}\times\tau=9$ used for embedding in Section 4.5 (recall that COT representations are obtained at every 10 steps). We also set $T_{f}$ to smaller values, e.g. $T_{f}=0.5$, and check that the qualitative results do not change.
In the following, we observe $a^{-1}\Delta\omega$ to check the relation (15). More specifically, we consider the average difference in the region $D_{1}$ by adding the forcing in the region $D_{0}$, which is denoted by
@@
Similarly, we define $\left\langle a^{-1}\Delta\omega\right\rangle_{D_{1}\to D_{0}}$. The average can be regarded as a suitable average of the linear response function $G_{\omega}({\bm{x}},t|{\bm{x}}^{\prime},t^{\prime})$ in view of equation (12). We also calculate the difference of the streamfunction in the same manner (we only consider the scalar quantities here).
@@
For $Re=16000$, we plot the averages of the differences in vorticity and stream function as a function of time in Figure 24. The differences are normalised by the forcing amplitude and the standard deviations of the observed regions. The amplitudes of the vorticity differences do not agree with our anticipation (14) or (15) as shown in Figure 24(a). On the other hand, for the streamfunction, with which we identified the structures $c^{1}_{+}$ and $c^{0}_{+}$, the amplitude of the difference for $D_{1}\to D_{0}$ is non-zero in $t\leq 15$, while the amplitude for $D_{0}\to D_{1}$ is around zero, as observed in Figure 24(b). However, in $t-t_{f}>15$ the amplitudes become comparable. From this behaviour, we infer that $c^{1}_{+}$ has a causal influence on $c^{0}_{+}$ up to around $t-t_{f}=T_{f}$ Baldovin et al. (2020) and that at later times no difference in causality is detected. Therefore, it is not straightforward to conclude which corner eddy is causal in another one by observing the difference of the amplitudes. For the lower Reynolds number $Re=14000$, where the flow is periodic in time, we plot the averaged differences as the linear response functions in Figure 25. Their overall shapes remain similar to those for $Re=16000$. Hence, we hardly find qualitative differences in linear responses across these Reynolds numbers. This may be owing to the linearity of their governing equations and comparable Péclet numbers despite the underlying flow transitioning from periodic to chaotic states. As a result, we find virtually no difference in sensitivity for periodic and chaotic evolutions. In this sense, the observational CCM approach based on TFDA can clarify the geometric causality between topological changes in flow patterns that the standard sensitivity analysis cannot capture. The discrepancy between CCM and the linear response analyses can be attributed to their fundamentally different approaches to causality detection. CCM operates in the framework of state-space reconstruction, utilising Taken’s embedding theorem to predict future states from historical time-series data. In the periodic regime illustrated in Figure 11(a), the COT representations exhibit highly regular patterns. Under these conditions, the system’s future states can be predicted from any instantaneous state with equal accuracy, regardless of which corner’s information is used. This makes causal relationships appear symmetric in the CCM analysis. Linear response analysis, in contrast, provides a mechanistic view of causality by examining the system’s response to controlled perturbations. This approach directly probes the spatio-temporal propagation of disturbances through the flow field. This methodological distinction highlights the complementary nature of observational (CCM) and interventional (linear response) approaches in understanding complex fluid systems.
@@
Finally, we discuss the relations of our work with recent studies on causality in fluid mechanics. In three-dimensional turbulent flows, a standard way is to apply POD first and then study the causal relationships between the POD modes. In Lopez-Doriga et al. (2024), the square duct flow was studied to uncover the cause of the secondary flow with the POD and the Granger causality analysis. In Martínez-Sánchez et al. (2023), the wake behind a circular cylinder mounted on a wall was analysed with the POD and the entropy-transfer method for causality. A time-series-based analysis for sub-grid energy fluxes with different grid scales in homogeneous and isotropic turbulence was performed in Lozano-Durán and Arranz (2022) with the entropy-transfer method. An elaborate interventional study to date in this field is Lozano-Durán et al. (2021) in which wall-bounded minimal-flow unit turbulence was analysed by numerically removing certain unstable modes to find the primary cause of the sustainment of turbulent fluctuations by the mean flow. In this paper, we studied the causality between the corner eddies represented by the COT symbols $c^{0}_{+}$ and $c^{1}_{+}$ as discrete modes in the lid-driven cavity flow. The difference from the previous work is that we used new modes and applied both approaches to causal analysis. The advantages of the new modes, COTs, are the immediate interpretability and the robustness to noise owing to the topological approach. The COTs are not much influenced by noise, unlike the POD. The disadvantage is that it does not apply directly to three-dimensional flows, but we need to consider the projection of the flow onto an appropriate two-dimensional section, as we discussed in Summary.
Appendix B From Hamiltonian to COT representation
This appendix outlines the procedure for obtaining the COT representation from a given Hamiltonian. The process involves two main steps: first, the construction of a Reeb graph from the Hamiltonian scalar field and second, the conversion of this Reeb graph into its corresponding COT representation.
B.1. Constructing the Reeb Graph from a Hamiltonian
Let $H(x,y)$ be the Hamiltonian for a given 2D vector field. The flow streamlines are the level sets of this function. The Reeb graph is a skeletal representation that captures the topological evolution of these level sets. Each point in a Reeb graph corresponds to a single connected component of a level-set of $H$. As the value of level $c$ is varied, the topology of the level set $H(x,y)=c$ changes: components may be created, destroyed, merged, or split. These topological events occur at the critical points of the Hamiltonian, which correspond to critical points in the flow field such as saddles and centres. The critical points of $H$ form the vertices of the Reeb graph, and the edges represent the evolution of the level sets as $c$ varies.
  • The local minima (births) and maxima (deaths) of $H$ correspond to the leaf vertices, where connected components of the level sets appear or vanish.
  • Saddle points, where level-set components merge or split, become vertices of a degree greater than two.
  • Edges represent continuous changes in topology between critical levels and can be weighted by the difference in $H$-values between endpoints.
For discrete data (e.g., obtained from simulation or PIV), calculating analytically critical points is not feasible. Instead, “birth” and “death” of connected components in sublevel sets ($H\leq c$) and super-level sets ($H\geq c$) are tracked by sweeping the Hamiltonian values from minimum to maximum. This generates two trees: a merge tree for sublevel sets and a split tree for super-level sets. These two trees are then merged to form the full Reeb graph. Note that the Reeb graph is equivalent to the COT by construction Uda et al. (2019).
B.2. Converting Reeb Graph to COT representation
The conversion from a Reeb graph to a COT representation is performed recursively.
  1. (1)
    When a threshold $\varepsilon>0$ is specified, the Reeb graph is simplified by removing edges with weights smaller than this threshold.
  2. (2)
    The traversal starts from a designated root of the Reeb graph. For a flow in a bounded domain, a boundary component with the smallest value of $H$ is typically chosen as the root and given the symbol $\beta_{\emptyset}$.
  3. (3)
    The algorithm traverses the graph from the root. At each vertex in the Reeb graph, it identifies the local structure on the basis of the vertex’s degree and the nature of the corresponding critical point. A COT symbol is assigned on the basis of the local structure.
    • Degree-1 vertices map to elliptic centres ($\sigma_{\pm}$).
    • Degree-3 vertices with self-connections map to $b$-type symbols ($b_{\pm\pm}$ or $b_{\pm\mp}$).
    • Degree-($2+n$) vertices connecting distinct saddles on the boundary map to $c_{\pm}$, where $n$ is the number of local orbit structures which are attached to the boundary and are enclosed by $c_{\pm}$ together with a part of the boundary.
    • The specific subscripts (++, -) are determined by the direction of the flow in the corresponding regions, which can be inferred from the Hamiltonian’s values.
  4. (4)
    For each incident edge (in cyclic order), it recurs on the child subgraph, obtaining its COT symbols. These symbols become arguments of the parent COT symbol in the same cyclic order.
The final COT string is constructed by concatenating the symbols assigned at each vertex, respecting the recursive structure of the Reeb graph.